# CodeRAG-Bench: Can Retrieval Augment Code Generation?

### v1

@@ 3.2 Retrieval Results
Table 3 shows diverse retrieval models’ performance across six tasks.
Comparison of Lexical and Neural Retrievers  BM25 has been widely used as a primary retrieval model in recent RACG work [47, 15], yet comprehensive comparisons against diverse retrieval systems are often under-explored. While prior studies indicate that neural retrieval systems often underperform BM25 baselines in out-of-domain scenarios [39], our analysis of CodeRAG-Bench reveals that dense embedding models frequently surpass BM25. We hypothesize that this is because many competitive retrieval models are trained on diverse tasks across various domains, including code data  [1, 36], enhancing their robustness in code retrieval setups.
@@
Efficiency  While larger retrieval models often outperform smaller ones, they often introduce significant costs. We analyze efficiency, focusing on (i) encoding latency: latency to encode documents offline, and (ii) search latency: latency to encode queries/documents and calculate their similarities, (iii) model storage requirements, and (iv) index storage requirements. We conduct efficiency analysis on sampled CodeSearchNet Python data. [^9] Due to the costs, we randomly sample 10k queries and 100k from CodeSearchNet Python split. For API models, we use a batch size of 64 for encoding. See experimental details in §C. As shown in Table 4,
Table 4: Efficiency analysis for document retrieval.
@@
BM25 indexing and searching takes only seconds to finish. Compared to base-size GIST-base, the SFR-Mistral model is more powerful in retrieval, yet requires over 5$\times$ larger index storage, and adds nearly 100$\times$ latency to encode documents, suggesting that the efficiency aspect should also be carefully studied for RAG pipelines.
3.3 Generation with and without Canonical Documents
We first evaluate possible lower- and upper-bounds on RACG performance by testing generation (i) without any retrieval, and (ii) with ground-truth documents. We report both results in Table 5. Compared to the base generation without contexts, incorporating canonical contexts improves in most setups, and substantially so on basic programming problems.
Table 5: Code generation pass@1 (i) without additional contexts (before the slash), and (ii) with ground-truth documents (after the slash). We only report (i) for LCB because LCB does not have ground-truth documents. We highlight results showing (ii) > (i) with green (darker green when having 10+ increases), and otherwise with red.
@@ 3.3 Generation with and without Canonical Documents
HumanEval MBPP LCB DS-1000 ODEX RepoEval SWE-bench
(function) (Lite)
StarCoder2-7B 31.7 / 94.5 10.4 / 34.8   1.5 29.2 / 30.0 14.6 / 17.5 26.5 / 42.0 0.0 / 0.7
CodeGemma-7B 49.4 / 77.4 48.0 / 52.2 21.5 20.1 / 19.8 18.9 / 18.2 24.7 / 32.2 0.0 / 0.3
CodeLlama-7B 34.8 / 87.2 23.8 / 42.8 13.5 21.8 / 26.1 35.8 / 41.0 24.1 / 38.3 0.3 / 0.0
CodeLlama-34B 42.7 / 84.8 51.2 / 88.0   5.8 34.7 / 37.0 34.9 / 38.0 29.8 / 42.6 0.0 / 0.0
DeepSeekCoder-7B 70.1 / 87.8 60.8 / 63.6 30.5 41.4 / 43.2 39.2 / 41.7 28.2 / 43.7 0.7 / 0.0
DeepSeekCoder-33B 78.0 / 95.7 61.0 / 92.2 33.8 40.2 / 40.1 28.0 / 28.9 32.4 / 45.3 0.3 / 0.7
Llama3-8B 57.9 / 65.2 35.6 / 52.8   2.8 28.9 / 31.1 37.4 / 33.7 26.0 / 43.2 0.7 / 0.3
Command-R 43.3 / 51.2 37.2 / 37.8 10.0 25.8 / 28.5 35.5 / 36.0 23.9 / 37.0 0.0 / 0.3
GPT-3.5-turbo 72.6 / 91.5 70.8 / 72.6 35.3 43.7 / 42.9 41.7 / 40.3 23.9 / 39.1 0.7 / 2.7
GPT-4 75.6 / 92.6 79.4 / 81.4 43.8 52.7 / 51.2 44.6 / 44.2 32.4 / 46.1 2.3 / 2.3
On open-domain problems, most code-specific LMs experience increases of up to $5.2$ points, signifying that most models can effectively consume indirectly helpful documentation. Among all general LMs, Command-R appears to be the only model benefiting from contexts for both datasets, consistent with its superior RAG ability. However, the strongest GPT does not gain from contexts, likely due to its prior familiarity with these libraries, as well-trained models are known to memorize popular facts and benefit little from retrieval for them [25, 16].
On repository-level problems, all models increase by 7.57.57.57.517.217.217.217.2 points from canonical snippets in RepoEval. SWE-bench Lite, however, is much harder, with even the strongest GPT model achieving only $2.7\%$ in the canonical setting. We conjecture that the difficulty comes from both complex multi-file editing and long contexts that stress the limits of most models. We leave the endeavor to integrate better inference-time strategies [45] with RAG to future work.
Larger models gain as much as their smaller variants, showing that RACG generally helps models of varied capacities. However, to unleash the full power of RACG, weaker LMs need to (i) extend their context lengths to take more documents, and (ii) further enhance their ability to consume contexts.
@@ 3.4 Retrieval-Augmented Code Generation
We now experiment with top-performing retrieval and generation models in the full RACG setting, which requires both retrieve documents and generating conditioned on the documents. We select the best retrieval models from each type: BM25, GIST-large, and OpenAI and Voyager embeddings. For generation, we select (i) StarCoder2-7B: a weaker model that benefits the most from contexts; (ii) DeepSeekCoder-7B: one of the strongest open code LMs; and (iii) GPT-3.5-turbo: a top proprietary model. [^10] We use deepseek-coder-7b/gpt-3.5-turbo instead of deepseek-coder-33b/gpt-4 due to resource limitations. For each dataset, we retrieve the most relevant contexts from its canonical source marked in Table 1, [^11] For HumanEval and MBPP, we exclude the canonical document for each query and retrieve top 5 documents. and retrieve programming solutions for LiveCodeBench. Table 6 shows the results.
Table 6: Performance of retrieval-augmented code generation, with top-performing retrieval and generation models. We bold-type the best RACG result for each dataset. LCB stands for LiveCodeBench.
Method General Open-Domain Repo-Level
HumanEval MBPP LCB DS-1000 ODEX RepoEval SWE-bench-Lite
@@
None 31.7 2.4 1.5 29.2 14.6 26.5 0.0
BM25 43.9 51.8 1.0 36.7 14.1 36.7 0.0
GIST-large 38.7 50.4 0.5 35.9 17.3 40.8 0.3
Voyage, code 39.0 52.6 0.3 36.0 15.3 45.8 0.3
OpenAI, small 39.0 52.6 1.5 35.5 15.9 51.2 0.0
OpenAI, rerank 34.8 53.4 0.5 33.4 14.1 53.9 0.3
Gold 94.5 34.8 - 30.0 17.5 42.0 0.7
@@
None 70.1 60.8 30.5 41.4 39.2 28.2 0.7
BM25 68.9 60.0 31.8 36.6 37.8 37.3 0.0
GIST-large 66.3 56.6 33.8 35.9 34.9 44.5 0.3
Voyage, code 66.5 56.4 31.8 35.9 39.4 46.6 0.3
OpenAI, small 68.9 58.6 32.0 35.5 37.1 55.2 0.3
OpenAI, rerank 53.0 60.6 31.5 36.5 37.1 55.5 0.3
Gold 87.8 63.6 - 43.2 41.7 48.1 0.0
w/ GPT-3.5-turbo
None 72.6 70.8 35.3 43.7 41.7 23.9 0.7
BM25 73.2 72.4 35.5 36.9 41.0 30.8 1.0
GIST-large 73.2 68.2 34.8 36.7 36.2 38.3 0.3
Voyage, code 75.0 66.8 34.5 37.4 41.0 43.2 0.7
OpenAI, small 73.8 68.4 35.8 36.9 40.3 48.0 0.3
OpenAI, rerank 64.0 72.6 33.5 37.4 40.5 49.6 0.3
Gold 91.5 72.6 - 42.9 40.3 39.1 2.7
Basic programming problems  Most retrieved contexts can help StarCoder2 generations. On MBPP, RACG even outperforms canonical setup by 15.615.615.615.617.817.817.817.8. However, RACG does not improve DeepSeekCoder generations, which we observe is due to over-complicated and ungrammatically repetitive generations when having additional contexts. This may indicate that DeepSeekCoder is not robust enough to extra contexts, and hence produces undesired behaviors when receiving different inputs. In comparison, GPT-3.5-turbo can effectively improve with added contexts, showing its better ability to leverage augmented contexts.
Open-domain problems  StarCoder2 substantially benefits from retrieved library documentation on both datasets, while DeepSeekCoder only improves on ODEX, and GPT-3.5 on neither. We hypothesize that the less familiar the model is with the domain, the more the model benefits from retrieving documents. Meanwhile, poor retrieval results can also impair the effectiveness of RACG. For example, DeepSeekCoder benefits from ground-truth documents (Sec. 3.3), but not from presumably lower-quality model-retrieved documents, indicating the need for better code retrieval models.
Repository-level problems  All models can benefit from retrieved code snippets on RepoEval, and RACG with openai-embeddings can often surpass the canonical setup. While some files do not include solutions (as in the canonical documents) to the problem, they may contain function definitions or usage examples that benefit end generation, suggesting that openai-embeddings understands the repository well and thus is able to retrieve implicitly supporting contexts. However, SWE-bench-lite is too complex and no RACG setups can get a non-trivial result.
3.5 How Many Documents to Augment?
[…]
@@ 4 RACG with Open Retrieval
Besides retrieving documents from the canonical source, we explore RACG with open retrieval from all sources (§2.2). We experiment with three category-representative datasets (HumanEval, ODEX, and RepoEval) as in §3.5. We also experiment with mixed retrieval documents from all sources, where we aggregate the top-1 documents from all five sources as additional contexts. [^12] For all experiments in this section, we only include the first 500 tokens of each retrieved document, which we show to be optimal on average in ablation studies in §4, and satisfies the context limits of all models.
4.1 Can RACG Benefit Weaker Models?
We use the three top-performing retrievers and the StarCoder2 generation model, as in §3.4, to examine if RACG helps weaker code LMs.
Table 7: Comparing five retrieval sources on HumanEval. We highlight results better than the no-retrieval baseline 31.7 with green, bold-type the best results for each source, and mark results with the canonical source in gray.
Method Program Tutorial Docs SO GitHub All
BM25 97.6 27.4 29.3 32.9 30.5 97.6
GIST-large 67.1 34.8 26.7 32.3 32.9 69.1
OpenAI 97.6 29.3 24.4 36.0 31.1 97.6
General Programming: HumanEval  Among all sources, SO posts can improve the results by 1.81.81.81.84.34.34.34.3, regardless of the choice of retrievers. Tutorials can improve results by $2.1$ only with the GIST-large neural retriever, possibly due to its adaptability. From manual examinations of the results, many retrieved posts and tutorials are about the same programming problem as the HumanEval example, with code and detailed textual explanations, hence could hint or disclose the answer. Other retrieval sources do not often contain relevant contexts thus do not improve generation. Surprisingly, generation with mixed documents performs as well as using the gold documents, suggesting that the model can discern and integrate the most useful content from a mixture of texts.
Table 8: Open retrieval on ODEX; no-retrieval baseline 14.6.
Method Program Tutorial Docs SO GitHub All
BM25 18.2 13.4 14.1 11.6 15.9 16.2
GIST-large 14.6 15.7 17.3 11.4 15.5 17.1
OpenAI 18.7 14.1 15.9 10.9 16.9 15.3
Open-Domain: ODEX  Programming solutions are the most helpful source, bringing gains of 3.83.83.83.84.34.34.34.3; GitHub files also improve by 0.90.90.90.92.32.32.32.3 points. Although the retrieved solutions/files are only sometimes functionally relevant to the ODEX examples, they can demonstrate the correct usage of libraries such as regex from solutions and requests from GitHub files, thus guiding the generation to be more functionally correct.
Similar to HumanEval, GIST-large is particularly good at retrieving tutorials, while BM25 and OpenAI embeddings find higher-quality program solutions, indicating their respective domain advantages.
Table 9: RACG with open retrieval on RepoEval; no-retrieval baseline 26.5.
Method Local Program Tutorial Docs SO GitHub Open L+O
BM25 36.7 23.6 25.2 23.9 23.6 25.5 23.6 31.4
GIST-large 40.8 24.1 23.3 21.7 24.7 24.4 24.1 41.8
@@ 4.1 Can RACG Benefit Weaker Models?
Repository-Level: RepoEval  Open sources are less useful than using code snippets retrieved from the local repository. As the RepoEval task is code completion, it is crucial to understand the local code context, which cannot be obtained from external sources. When using both local and open-source contexts (L+O), models surpass the no-retrieval baseline, yet are still only comparable with Local, suggesting more efforts to build systems that benefit from both sources.
Figure 3: Performance variation with different chunking sizes.
Exploring optimal chunking strategies  Adding multiple documents may exceed model context limits hence imparing RACG. Therefore, we explore various chunking strategies to better include retrieved contexts. Compared to the no-chunking baseline, we study (i) post-retrieval chunking that takes the first N-tokens of each document, (ii) post-retrieval with reranking using BGE-reranker-base (§3.1) to find the most relevant N-token chunk from each document, and (iii) pre-retrieval chunking that chunks documents beforehand and retrieves N-token pieces directly. [^13] We do not chunk programming solutions since they are typically short (average ¡200 tokens as in Table 2).
For (i), we compare using the first N-tokens for N from 200 to 1500. As in Figure 3, most sources are best represented by the first 800 tokens; SO posts perform best with the first 200 tokens. We then perform (ii) reranking within this optimal range of 200–800 tokens, yet find it greatly degrades the results, showing the limited utility of current rerankers. Lastly, Table 10 shows that (iii) pre-retrieval achieves the highest scores on almost all document sources.
Table 10: Comparing different chunking strategies on HumanEval with $N=500$.
@@
4.2 Does RACG Help Stronger Models?

### v2

@@ 3.2 Retrieval Results
Table 3 shows retrieval results on six tasks.
Comparison of lexical and neural retrievers  BM25 has been widely used as a primary retrieval model in recent RACG work (Zhou et al., 2023; Jimenez et al., 2024), yet comprehensive comparisons against diverse retrieval systems are often under-explored. While prior studies indicate that neural retrieval systems often underperform BM25 baselines in out-of-domain scenarios (Thakur et al., 2021), our analysis of CodeRAG-Bench reveals that dense embedding models frequently surpass BM25. We hypothesize that this is because many competitive retrieval models are trained on diverse tasks across various domains, including code data  (Asai et al., 2023; Su et al., 2023), enhancing their robustness in code retrieval setups.
Do code retrieval models perform better? At similar parameter scales, models specifically trained for code retrieval tasks typically show superior performance. Notably, Jina-v2-code outperforms GIST-base and BGE-base by 7.4 and 6.6 average NDCG@10, respectively, while Voyage-code significantly outperforms OpenAI-03.
@@
Table 4: Efficiency analysis for document retrieval.
Efficiency  While larger retrieval models often outperform smaller ones, they often introduce significant costs. We analyze efficiency, focusing on (i) encoding latency: latency to encode documents offline, and (ii) search latency: latency to encode queries/documents and calculate their similarities, (iii) model storage requirements, and (iv) index storage requirements. We conduct efficiency analysis on sampled CodeSearchNet Python data. [^7] Due to the costs, we randomly sample 10k queries and 100k from CodeSearchNet Python split. For API models, we use a batch size of 64 for encoding. See experimental details in §B. As shown in Table 4, BM25 indexing and searching takes only seconds to finish. Compared to base-size GIST-base, the SFR-Mistral model is more powerful in retrieval, yet requires over 5$\times$ larger index storage, and adds nearly 100$\times$ latency to encode documents, suggesting that the efficiency aspect should also be carefully studied for RAG pipelines.
@@
HumanEval MBPP LCB DS-1000 ODEX ODEX-hard RepoEval SWE-bench
w/o gold w/o gold w/o w/o gold w/o gold w/o gold w/o gold w/o gold
StarCoder2-7B 31.7 94.5 10.4 34.8   1.5 29.2 30.0 14.6 17.5 10.3 17.2 26.5 42.0 0.0   0.7
CodeGemma-7B 49.4 77.4 48.0 52.2 21.5 20.1 19.8 18.9 18.2 13.8 13.8 24.7 32.2 0.0   0.3
CodeLlama-7B 34.8 87.2 23.8 42.8 13.5 21.8 26.1 35.8 41.0 27.6 31.0 24.1 38.3 0.0   0.0
CodeLlama-34B 42.7 84.8 51.2 88.0   5.8 34.7 37.0 34.9 38.0 17.2 27.6 29.8 42.6 0.0   0.0
DeepSeekCoder-7B 70.1 87.8 60.8 63.6 30.5 41.4 43.2 39.2 41.7 17.2 24.1 28.2 43.7 0.0   0.0
DeepSeekCoder-33B 78.0 95.7 61.0 92.2 33.8 40.2 40.1 28.0 28.9 24.1 31.0 32.4 45.3 0.3   0.7
Llama3-8B 57.9 65.2 35.6 52.8   2.8 28.9 31.1 37.4 33.7 13.8 17.2 26.0 43.2 0.0   0.3
Command-R 43.3 51.2 37.2 37.8 10.0 25.8 28.5 35.5 36.0 20.7 20.7 23.9 37.0 0.0   0.3
GPT-3.5-turbo 72.6 91.5 70.8 72.6 35.3 43.7 42.9 41.7 40.3 17.2 24.1 23.9 39.1 0.7   6.3
GPT-4o 75.6 92.6 79.4 81.4 43.8 52.7 51.2 44.6 44.2 20.7 27.6 32.4 46.1 2.3 30.7
Table 5: Code generation pass@1 without additional contexts (w/o), and with ground-truth documents (gold). We only report w/o for LCB because LCB does not have ground-truth documents. We highlight results showing gold > w/o with green (darker green when having 10+ increases), and with red if gold < w/o.
3.3 Generation with Canonical Documents
We first evaluate possible lower- and upper-bounds on RACG results by testing generation (i) without any retrieval, and (ii) with ground-truth documents. We report both results in Table 5. Compared to the base generation without contexts, incorporating canonical contexts improves in most setups, and substantially so on basic programming problems.
On open-domain tasks, most code-specific LMs increase up to $5.2$ points, signifying that most models can benefit from indirectly helpful documents. In contrast, GPTs show no gains with retrieval. We hypothesize that this is because both datasets mostly test on common Python libraries, which powerful models may have already memorized, similar to their memorization of factual knowledge (Mallen et al., 2023; Kandpal et al., 2023), thereby reducing the need for retrieval. To verify this hypothesis, we build an ODEX subset of examples with the 20 least used libraries, i.e., ODEX-hard. As shown in Table 5, adding documents retrieved with most methods improves the results by 20.3–40.1%, showing the effectiveness of RACG on challenging coding tasks using unfamiliar libraries.
Repository-level challenges All models show gains of 7.5–17.2 points with canonical snippets in RepoEval, but SWE-bench Lite proves much more challenging — only GPT-3.5-turbo and GPT-4o achieve non-trivial results, consistent with previous findings (Yang et al., 2024). Notably, GPT-4o shows a 27.4% increase when using gold documents on SWE-bench, indicating that retrieval significantly enhances performance when paired with strong core generation capabilities, even in highly challenging coding tasks.
@@ 3.4 Retrieval-Augmented Code Generation
We now experiment with top-performing retrieval and generation models in the full RACG setting, which requires both retrieve documents and generating conditioned on the documents. We select the best retrieval models from each type: BM25, GIST-large, Voyage, and OpenAI embeddings. For generation, we select (i) StarCoder2-7B: a weaker model that benefits the most from contexts; (ii) DeepSeekCoder-7B: one of the strongest open code LMs; and (iii) GPT-3.5-turbo: one of the top proprietary models. For each dataset, we retrieve the most relevant contexts from its canonical source marked in Table 1, and retrieve programming solutions for LiveCodeBench. Table 6 shows the results. Note that we exclude canonical docs (answers) from the retrieval corpora for basic programming tasks.
Overall, the best retrieval models vary depending on the task and underlying LMs. In some cases, top-performing retrieval models do not lead to the best RACG outcomes, highlighting the need to evaluate RACG systems holistically across varied tasks.
Method Basic Programming Open-Domain Repo-Level
HumanEval MBPP LCB DS-1000 ODEX ODEX-hard RepoEval SWE-bench
@@
None 31.7 2.4 1.5 29.2 14.6 10.3 26.5 0.0
BM25 43.9 51.8 1.0 36.7 14.1 13.8 36.7 0.0
GIST-large 38.7 50.4 0.5 35.9 17.3 13.8 40.8 0.3
Voyage, code 39.0 52.6 0.3 36.0 15.3 10.3 45.8 0.3
OpenAI, small 39.0 52.6 1.5 35.5 15.9 17.2 51.2 0.0
OpenAI, rerank 34.8 53.4 0.5 33.4 14.1 17.2 53.9 0.3
Gold 94.5 34.8 - 30.0 17.5 17.2 42.0 0.7
@@
None 70.1 60.8 30.5 41.4 39.2 17.2 28.2 0.7
BM25 68.9 60.0 31.8 36.6 37.8 20.7 37.3 0.0
GIST-large 66.3 56.6 33.8 35.9 34.9 20.7 44.5 0.3
Voyage, code 66.5 56.4 31.8 35.9 39.4 17.2 46.6 0.3
OpenAI, small 68.9 58.6 32.0 35.5 37.1 20.7 55.2 0.3
OpenAI, rerank 53.0 60.6 31.5 36.5 37.1 24.1 55.5 0.3
Gold 87.8 63.6 - 43.2 41.7 24.1 48.1 0.0
w/ GPT-3.5-turbo GPT-4o
None 72.6 70.8 35.3 43.7 41.7 17.2 23.9   2.3
BM25 73.2 72.4 35.5 36.9 41.0 24.1 30.8   6.7
GIST-large 73.2 68.2 34.8 36.7 36.2 13.8 38.3 19.3
Voyage, code 75.0 66.8 34.5 37.4 41.0 20.7 43.2 15.7
OpenAI, small 73.8 68.4 35.8 36.9 40.3 17.2 48.0 21.0
OpenAI, rerank 64.0 72.6 33.5 37.4 40.5 17.2 49.6 21.7
Gold 91.5 72.6 - 42.9 40.3 24.1 39.1 30.7
Table 6: Performance of retrieval-augmented code generation, with top retrieval and generation models. We bold-type the best RACG results. We test gpt-4o on SWE-bench to show non-trivial results than gpt-3.5-turbo. Note that we exclude code canonical answer from the retrieval corpora for basic programming tasks.
Basic programming problems  Most retrieved contexts help StarCoder2 generations. On MBPP, RACG even outperforms canonical setup by 15.615.615.615.617.817.817.817.8. However, RACG does not improve DeepSeekCoder generations, which we observe is due to over-complicated and ungrammatically repetitive generations when with additional contexts. In comparison, GPT-3.5-turbo can effectively improve with added contexts, showing its better ability to leverage augmented contexts.
Open-domain problems  The weaker StarCoder2 benefits from retrieved library documentation across all datasets, while DeepSeekCoder and GPT-3.5 show gains mainly on ODEX-hard problems. This aligns with findings from the canonical document setup, indicating that RACG is particularly effective for less popular libraries. Interestingly, despite relatively low NDCG@10 scores, the best-performing RACG combinations match their canonical results on ODEX-hard.
Repository-level problems  All models show improvements with retrieved code snippets on RepoEval, with RACG using strong retrievers like openai-embeddings performing on par with—or even surpassing—the canonical setup, likely due to the additional context provided to the models. On SWE-Bench, the best-performing combination, Retrieval-then-Rerank and GPT4o, yields a 21-point improvement over the no-retrieval baseline. However, there remains a 9-point gap compared to the gold setup, indicating room for improvement on the retrieval side, as reflected in the limited code retrieval performance shown in Table 3.
Model Retriever HumanEval ODEX
w/o Prog Tut Docs SO GitHub All w/o Prog Tut Docs SO GitHub All
StarCoder BM25 31.7 97.6 27.4 29.3 32.9 30.5 97.6 14.6 18.2 13.4 14.1 11.6 15.9 16.2
GIST 67.1 34.8 26.7 32.3 32.9 69.1 14.6 15.7 17.3 11.4 15.5 17.1
OpenAI 97.6 29.3 24.4 36.0 31.1 97.6 18.7 14.1 15.9 10.9 16.9 15.3
GPT-4o OpenAI 75.6 94.5 90.2 90.9 91.5 84.8 95.1 44.6 49.2 44.2 47.6 40.3 39.4 39.6
Table 7: Comparing five retrieval sources on HumanEval and ODEX, using StarCoder2 (top) and GPT-4o (bottom).
@@ 4 RACG with Open Retrieval
Besides retrieving documents from the canonical source, we explore RACG with open retrieval from all sources (§2.2) on three category-representative datasets: HumanEval, ODEX, and RepoEval. We also study mixed retrieval with documents from all sources, where we aggregate the top-1 documents from all five sources as augmented contexts. [^8] We use the first 500 tokens of each document for all experiments in this section, which we show to be optimal in ablation studies (§4), and satisfies all model context limits. We use the three top retrievers along with StarCoder2 and OpenAI retrieval with GPT-4o generation, to study open RACG with weak and strong LMs.
General programming: HumanEval  Among all sources, SO posts can improve the results by 1.81.81.81.84.34.34.34.3, regardless of the choice of retrievers. Tutorials can improve results by $2.1$ only with the GIST retriever.From manual examinations of the results, many retrieved posts and tutorials are about the same programming problem as the HumanEval example, with code and detailed textual explanations, hence could hint or disclose the answer. Other retrieval sources do not often contain relevant content thus do not bring improvements. Surprisingly, generation with mixed documents performs as well as using the gold documents, suggesting that the model can discern and integrate the most useful content from a mixture of texts.
Open-domain: ODEX  Programming solutions are the most helpful source by bringing 3.8–4.3 gains, even surpassing gains of canonical documentation. Notably, both GPT-4o and StarCoder using OpenAI retrieval from programming solutions, outperform their variants retrieving from documentation by 3.2 and 1.6 points. Although the retrieved content is only sometimes functionally relevant to the ODEX examples, they can exemplify the correct usage of libraries such as regex in solutions and requests in GitHub files, thus guiding the generation to be more functionally correct. Similar to HumanEval, GIST-large is particularly good at retrieving tutorials, while BM25 and OpenAI embeddings find higher-quality program solutions, indicating their respective domain advantages.
Repository-level: RepoEval  Open sources are less useful than code snippets in the local repository. Understanding local code contexts is crucial and irreplaceable than external resources. When using both local and open-source contexts (L+O), models surpass the no-retrieval baseline, yet are still only comparable with Local, suggesting more efforts and insights to benefit from both sources.
Method w/o Local Prog Tut Docs SO GitHub Open L+O
StarCoder2-7B
BM25 26.5 36.7 23.6 25.2 23.9 23.6 25.5 23.6 31.4
GIST 40.8 24.1 23.3 21.7 24.7 24.4 24.1 41.8
@@
GPT-4o
OpenAI 32.4 62.2 35.4 28.7 27.8 29.0 28.2 30.3 54.2
Table 8: RACG with open retrieval on RepoEval.
Exploring optimal chunking strategies  Adding many documents may exceed model context limits hence impairing RACG, we thus explore various chunking strategies to better integrate retrieval. Compared to the no-chunking baseline, we study (i) post-retrieval chunking that takes the first N-tokens of each document, (ii) post-retrieval with reranking using BGE-reranker-base (§3.1) to find the most relevant N-token chunk from each document, and (iii) pre-retrieval chunking that chunks documents beforehand and retrieves N-token pieces directly. [^9] We do not chunk programming solutions since they are typically short (average ¡200 tokens as in Table 2).
Figure 2: Performance with different chunking sizes.
We compare (i) using the first N-tokens for N from 200 to 1500 (Figure 2). Most sources are best represented by the first 800 tokens except for SO posts. However, we find (ii) reranking within this optimal range of 200–800 tokens greatly degrades the results, showing limited utility of current rerankers. Lastly, (iii) pre-retrieval achieves the highest scores on almost all document sources (Table 9).
@@
Table 9: Comparing chunking strategies on HumanEval.
5 Related Work
Code generation  Neural code generation has been a crucial task (Lu et al., 2021), and increasingly strong code LMs have been created (Roziere et al., 2023; Li et al., 2023; Guo et al., 2024; Team, 2024) to solve various tasks (Chen et al., 2021; Lai et al., 2023; Jimenez et al., 2024). However, most LMs generate code solely based on NL queries and model parametric knowledge, without using external programming sources (e.g., tutorials). To fill in this gap and allow systematic studies of RACG, we integrate various datasets and retrieval sources to build CodeRAG-Bench.
Retrieval augmented generation (RAG)  RAG has been widely used in knowledge-intensive tasks (Lewis et al., 2020; Guu et al., 2020), however, mostly on text-centric tasks using general domain corpora such as Wikipedia Asai et al. (2024). Some works used programming context retrieved from repositories Ding et al. (2023); Yang et al. (2024) or documentations Zhou et al. (2023), yet none of them considered RACG across varied coding tasks and knowledge sources. In text-centric tasks, unified benchmarks such as BEIR (Thakur et al., 2021) and KILT (Petroni et al., 2020) aggregate retrieval and generation tasks and facilitate its progress Muennighoff et al. (2022). To similarly enable systematic studies of RACG across coding tasks and retrieval sources, we curate a unified benchmark and release its RACG codebase.
6 Conclusion
In this work, we propose CodeRAG-Bench, a benchmark for retrieval-augmented code generation with various coding tasks and retrieval sources. With our experiments with top-performing retrieval and generation models, we show that retrieving external documents can greatly benefit code generation. However, current retrieval models struggle to find useful documents, and generation models have limited context capacity and RAG abilities, both leading to suboptimal RACG results. We hope CodeRAG-Bench can serve as a solid testbed to advance future endeavors in this direction.
Limitations
We propose a new paradigm, retrieval-augmented code generation, equipped with a comprehensive benchmark CodeRAG-Bench. However, as an initial exploration in this field, our work could be extended in task and language diversity, as well as model and methodological improvements.
We aggregate various existing code generation tasks, but many interesting scenarios such as code debugging remain under-explored. Meanwhile, we focus on coding tasks using Python programming language, but extrapolating to other languages may bring additional challenges.
Meanwhile, for benchmarking purposes, we mostly experimented with vanilla retrieval, reranking, and generation methods, but better backbone models and advanced methods for each RACG component are yet fully explored. Our results may not represent all model behaviors, and we encourage future works to build methods that break certain limitations we observe in current systems.
Acknowledgment
We thank Shuyan Zhou and Xinran Zhao for the helpful discussions in the early stage of this project; Saujas Vaduguru, Jing Yu Koh, Alex Xie, and Andy Liu for providing valuable feedback for the draft.
Appendix A Example Illustrations
A.1 Example with Canonical Documents
A.2 RACG with Helpful and Distracting Documents
Appendix B Additional Details about Retrieval Efficiency
Appendix C Result Reproduction
StarCoder2  The StarCoder2 technical report (Lozhkov et al., 2024) reported results on the HumanEval, MBPP, and DS-1000 datasets. On HumanEval, our reproduced results (31.7) is slightly lower than their number (35.4), possibly because the original paper additionally input the test cases as additional information in the prompt, whereas in our basic NL-to-code setup, no test cases are provided. This additional information may cause their results to be higher.
On MBPP dataset, they adopt a subset of MBPP, i.e., 399 out of 427 examples that have additional test cases populated by Liu et al. (2023). In contrast, we evaluate on the entire dataset, which is likely to cause the variance in results.
CodeGemma  The CodeGemma technical report (Team, 2024) reported results on HumanEval and MBPP datasets, but does not provide any details about the instructions, few-shot examples, or other parts of the prompt that they use. We were able to roughly reproduce their reported results, but with 3-5 points less in pass@1.
CodeLlama  The CodeLlama technical report (Roziere et al., 2023) reports results on HumanEval and MBPP datasets. We were able to perfectly reproduce their results on the HumanEval dataset under the zero-shot setting. However, for MBPP experiments, they use 3-shot prompting, which could potentially explain that our zero-shot results are 4 points lower in pass@1.
DeepSeekCoder  The DeepSeekCoder technical report (Guo et al., 2024) reports results on HumanEval and MBPP for the 7B-instruct-v1.5 and the 33B-instruct models, the report additionally report DS1000 results for the 33B-instruct model. We could reproduce the original results on HumanEval and DS-1000, but got slightly worse results on MBPP because they used few-shot prompting, which should outperform our zero-shot method.
Appendix D Analysis on Open-Domain Coding Problems
In §3.4, providing the documentation of required libraries brings limited benefits, especially with strong proprietary models such as GPT and Gemini. While we hypothesize that these strong models are sufficiently familiar with the required libraries and in turn barely benefit from additional information about them, in this section, we quantitatively investigate this issue and verify its validity.
Concretely, we construct a subset of ODEX containing only examples with less common libraries. We use the real-world distribution of all libraries involved in ODEX and select examples that use the top 20 least common libraries (e.g., sqlite3, ftplib, flask). We then evaluate model performance on this subset and compare the results with and without documentation in model contexts.
With varied retrieval models  Aligning with §3.4, we examine the RACG results using documentation retrieved by different retrieval models. As shown in Table 10, augmenting documentation retrieved with most methods improves the results by 20.3–40.1%. Compared to the entire ODEX set where most queries require common libraries, this hard-library split more clearly demonstrates the effectiveness of augmenting library documentation. This result verifies our hypothesis that strong GPT models are familiar with most common libraries, and can only benefit from additional library information when harder libraries are required.
Model none BM25 GIST Voyage OpenAI Gold
GPT-3.5-turbo 17.2 24.1 13.8 20.7 17.2 24.1
GPT-4 20.7 24.1 17.2 27.6 24.1 27.6
Table 10: RACG results on the subset of ODEX examples using the least common libraries.
Appendix E How Many Documents to Augment?
[…]
Appendix F Does RACG Help Stronger Models?